# Fact-checking information generated by a large language model can decrease news discernment

### v3

@@ 4.1 Attitude towards AI
Veracity Opt in -- Opt out Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.027 0.010 10430 2.807 0.030 *
False 0.010 0.009 10430 1.074 1.000
@@
$\dagger$ Bonferroni’s method comparing a family of 6 estimates
@@
Veracity Opt in -- Opt out Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.040 0.011 10510 3.742 0.001 **
False 0.032 0.011 10510 2.996 0.016 *
@@
$\dagger$ Bonferroni’s method comparing a family of 6 estimates
@@ 4.2 Headline congruence
We now examine the potential moderating effects of headline congruence on participants’ belief in and intention to share them, shown in Figure S4. We model this relationship by including a three-way interaction between Condition, Veracity, and headline Congruence (Condition $\times$ Veracity $\times$ Congruence). The results related to belief and sharing intent can be found in Tables S23 and S24, respectively. We find no evidence of a significant three-way interaction between headline congruence in either group, suggesting that discernment is not altered by the effects of headline congruence.
Figure S4: Relationship between (a) belief in and (b) intent to share headlines and their congruency across all conditions. Congruency is grouped along the x-axis based on headline congruency.
Table S23: Ineffectiveness of LLM Fact Checks Coefficients (Congruence interaction; Belief Group; $F=528.27$, $R^{2}=0.19$, $P<0.001$)
@@
(Intercept) 0.600 0.051 11.724 $<0.001$**
Cond.(Forced) 0.029 0.031 0.943 0.346
Cond.(Optional) -$2.653\times 10^{-4}$ 0.031 -0.009 0.993
Veracity(True) 0.409 0.031 13.322 $<0.001$**
Congr.(Inc.) -0.065 0.014 -4.683 $<0.001$**
Age -0.006 0.001 -7.444 $<0.001$**
Education 0.007 0.006 1.180 0.238
Cond.(Forced):Veracity(True) -0.056 0.042 -1.344 0.179
@@
Cond.(Forced):Congr.(Inc.) 0.017 0.010 1.588 0.112
Cond.(Optional):Congr.(Inc.) 0.026 0.007 3.956 $<0.001$**
Veracity(True):Congr.(Inc.) -0.014 0.024 -0.578 0.563
Cond.(For.):Veracity(True):Congr.(Inc.) 0.018 0.023 0.769 0.442
Cond.(Opt.):Veracity(True):Congr.(Inc.) 0.002 0.027 0.081 0.935
@@
Table S24: Ineffectiveness of LLM Fact Checks Coefficients (Congruence interaction; Share Group; $F=293.05$, $R^{2}=0.11$, $P<0.001$)
@@
(Intercept) 0.886 0.046 19.123 $<0.001$**
Cond.(Forced) 0.023 0.032 0.740 0.459
Cond.(Optional) -0.038 0.034 -1.131 0.258
Veracity(True) 0.102 0.026 3.912 $<0.001$**
Congr.(Inc.) -0.059 0.013 -4.747 $<0.001$**
Age -0.008 0.001 -10.418 $<0.001$**
Education -0.025 0.009 -2.788 0.005 **
Cond.(Forced):Veracity(True) -0.022 0.024 -0.923 0.356
@@
Cond.(Forced):Congr.(Inc.) 0.015 0.016 0.973 0.331
Cond.(Optional):Congr.(Inc.) 0.051 0.013 4.020 $<0.001$**
@@
Cond.(Forced):Veracity(True):Congr.(Inc.) 0.031 0.023 1.345 0.179
Cond.(Optional):Veracity(True):Congr.(Inc.) -0.002 0.012 -0.189 0.850
@@
Figure S5 illustrates the relationship between belief in headlines and their congruency across all fact-checking scenarios and experimental conditions. The same relationship is presented with respect to sharing intent in Figure S6. We model this relationship using a three-way interaction between Condition, fact-checking scenario, and headline congruence (Condition $\times$ FC Scenario $\times$ Congruence). Again, we focus on the forced and control conditions and exclude data for the optional participants when fitting each model. The results of fitting the belief and share group models are found in Tables S25 and S26. However, we again must utilize these models for post-hoc comparisons similar to those presented in the main text for each group. To do so, we recode congruent headlines as 0 and incongruent headlines as 1. Using this dummy coding, we then compare headline congruence fitted slopes between the Control and Forced groups. These results are shown in Tables S27 and S28 for the belief and share group, respectively, finding no evidence of significant interactions for either group in any FC scenario.
@@
Cond.(Forced):FC Scen.(True $\times$ false) -0.190 0.052 -3.683 $<0.001$**
@@
FC Scen.(False $\times$ unsure):Congr.(Inc.) -0.102 0.040 -2.537 0.011 *
@@
Cond.(Forced):FC Scen.(False $\times$ unsure):Congr.(Inc.) 0.008 0.045 0.173 0.862
@@
Cond.(Forced):FC Scen.(True $\times$ true):Congr.(Inc.) 0.069 0.035 1.962 0.050
@@
Cond.(Forced):FC Scen.(False $\times$ unsure):Congr.(Inc.) -0.088 0.056 -1.571 0.117
@@
Cond.(Forced):FC Scen.(True $\times$ true):Congr.(Inc.) 0.064 0.075 -0.855 0.393
@@
Headline Scenario Control -- Forced Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false -0.017 0.019 18738 -0.873 1.000
False $\times$ unsure -0.025 0.058 18738 -0.425 1.000
True $\times$ false -0.078 0.041 18738 -1.887 1.000
True $\times$ true -0.086 0.048 18738 -1.804 1.000
True $\times$ unsure -0.012 0.023 18738 -0.511 1.000
@@
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
@@
Headline Scenario Control -- Forced Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false -0.025 0.020 19938 -1.278 1.000
False $\times$ unsure 0.062 0.059 19938 1.055 1.000
True $\times$ false -0.038 0.042 19938 -0.919 1.000
True $\times$ true 0.039 0.048 19938 0.815 1.000
True $\times$ unsure -0.071 0.023 19938 -3.085 0.091 $\cdot$
@@
$\dagger$ Bonferroni’s method comparing a family of 45 estimates
Next, we examine whether behaviors in the optional condition differ based on the congruence of headlines by introducing a three-way interaction term involving whether a participant chose to view LLM fact checks (opt in vs. opt out), headline veracity, and headline congruence (Opt-Condition $\times$ Veracity $\times$ Congruence). Tables S29 and S30 show the results of fitting these models for the belief and intent to share groups, respectively. We perform a post-hoc comparison of the belief (Table S31) and sharing (Table S32) group slopes, fit to the opt-in and opt-out conditions across different levels of congruence. The results of these post-hoc comparisons are shown in Tables S33 and S34, respectively.
@@
[…]
Table S29: Opt In versus Opt Out Coefficients (Congruency interaction, Belief Group; $F=308.07$, $R^{2}=0.22$, $P<0.001$)
@@
(Intercept) 0.643 0.062 10.372 $<0.001$**
Option(opt out) -0.215 0.045 -4.762 $<0.001$**
Veracity(True) 0.278 0.045 6.125 $<0.001$**
Congr.(Inc.) -0.002 0.021 -0.085 0.933
Age -0.004 0.001 -3.836 $<0.001$**
Education -0.008 0.009 -0.872 0.384
Option(opt out):Veracity(True) 0.242 0.052 4.620 $<0.001$**
Option(opt out):Congr.(Inc.) -0.066 0.040 -1.659 0.097 $\cdot$
Veracity(True):Congr.(Inc.) -0.037 0.031 -1.202 0.229
Option(opt out):Veracity(True):Congr.(Inc.) 0.041 0.050 0.807 0.420
@@
Table S30: Opt In versus Opt Out Coefficients (Congruency interaction, Share Group; $F=260.78$, $R^{2}=0.19$, $P<0.001$)
@@
(Intercept) 0.796 0.068 11.793 $<0.001$**
Option(opt out) -0.301 0.038 -7.877 $<0.001$**
Veracity(True) 0.074 0.021 3.594 $<0.001$**
Congr.(Inc.) 0.015 0.006 2.485 0.013 *
Age -0.006 0.001 -4.981 $<0.001$**
Education -0.001 0.014 -0.054 0.957
Option(opt out):Veracity(True) 0.027 0.030 0.895 0.371
Option(opt out):Congr.(Inc.) -0.031 0.012 -2.496 0.013 *
Veracity(True):Congr.(Inc.) -0.034 0.012 -2.906 0.004 **
Option(opt out):Veracity(True):Congr.(Inc.) -0.012 0.026 -0.478 0.633
@@
Condition Veracity $b$ Std. Error df $t$-ratio $P$ Sig.
Opt in False -0.002 0.018 9990 -0.101 0.920
Opt out False -0.068 0.018 9990 -3.818 $<0.001$**
Opt in True -0.039 0.017 9990 -2.372 0.018 *
Opt out True -0.065 0.019 9990 -3.372 $<0.001$**
@@
Condition Veracity $b$ Std. Error df $t$-ratio $P$ Sig.
Opt in False 0.015 0.016 10150 0.938 0.348
Opt out False -0.016 0.019 10150 -0.829 0.407
Opt in True -0.019 0.016 10150 -1.183 0.237
Opt out True -0.062 0.020 10150 -3.064 0.002 **
@@
Veracity Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.026 0.025 9990 1.012 1.000
False 0.066 0.025 9990 2.642 0.050 *
@@
$\dagger$ Bonferroni’s method comparing a family of 6 estimates
@@
Veracity Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True 0.044 0.026 10150 1.699 0.536
False 0.031 0.025 10150 1.239 1.000

### v4

@@ 4.1 Attitude towards AI
Headline Scenario Opt in -- Opt out Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false 0.011 0.010 10418 1.142 1.000
False $\times$ unsure -0.003 0.030 10418 -0.110 1.000
True $\times$ false 0.029 0.021 10418 1.369 0.8550
True $\times$ true 0.004 0.024 10418 0.147 1.000
True $\times$ unsure 0.032 0.012 10418 2.678 0.037 *
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@
Headline Scenario Opt in -- Opt out Std. Err. df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false 0.029 0.011 10498 2.571 0.051 $\cdot$
False $\times$ unsure 0.059 0.033 10498 1.766 0.387
True $\times$ false 0.029 0.024 10498 1.217 1.000
True $\times$ true 0.040 0.027 10498 1.481 0.693
True $\times$ unsure 0.044 0.013 10498 3.315 0.005 **
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@ 4.2 Headline congruence
We now examine the potential moderating effects of headline congruence on participants’ belief in and intention to share them, shown in Figure S4. We model this relationship by including a three-way interaction between Condition, Veracity, and headline Congruence (Condition $\times$ Veracity $\times$ Congruence). The results related to belief and sharing intent can be found in Tables S25 and S26, respectively. We find no evidence of a significant three-way interaction between headline congruence in either group, suggesting that average discernment is not altered by the effects of headline congruence.
Figure S4: Relationship between (a) belief in and (b) intent to share headlines and their congruency across all conditions. Headline congruency is shown along the x-axis.
Table S25: Ineffectiveness of LLM Fact Checks Coefficients (Congruence interaction; Belief Group; $F=762.09$, $R^{2}=0.25$, $P<0.001$)
@@
(Intercept) 0.582 0.049 11.857 $<0.001$**
Cond.(Forced) 0.029 0.031 0.930 0.352
Cond.(Optional) -0.001 0.031 -0.029 0.977
Cond.(HumanFC) -0.030 0.030 -1.006 0.314
Veracity(True) 0.409 0.031 13.311 $<0.001$**
Congr.(Inc.) -0.065 0.014 -4.646 $<0.001$**
Age -0.006 0.001 -7.513 $<0.001$**
Education 0.009 0.005 1.874 0.061 $\cdot$
Cond.(Forced):Veracity(True) -0.056 0.042 -1.341 0.180
@@
Cond.(HumanFC):Veracity(True) 0.184 0.034 5.389 $<0.001$**
Cond.(Forced):Congr.(Inc.) 0.017 0.011 1.560 0.119
Cond.(Optional):Congr.(Inc.) 0.026 0.006 4.058 $<0.001$**
Cond.(HumanFC):Congr.(Inc.) 0.035 0.007 4.718 $<0.001$**
Veracity(True):Congr.(Inc.) -0.014 0.024 -0.573 0.567
Cond.(Forced):Veracity(True):Congr.(Inc.) 0.018 0.023 0.759 0.448
Cond.(Optional):Veracity(True):Congr.(Inc.) 0.002 0.027 0.082 0.935
Cond.(HumanFC):Veracity(True):Congr.(Inc.) -0.010 0.017 -0.620 0.535
@@
Table S26: Ineffectiveness of LLM Fact Checks Coefficients (Congruence interaction; Share Group; $F=313.41$, $R^{2}=0.11$, $P<0.001$)
@@
(Intercept) 0.879 0.044 20.031 $<0.001$**
Cond.(Forced) 0.023 0.032 0.722 0.470
Cond.(Optional) -0.041 0.034 -1.208 0.227
Cond.(HumanFC) -0.017 0.036 -0.460 0.646
Veracity(True) 0.102 0.026 3.913 $<0.001$**
Congr.(Inc.) -0.059 0.013 -4.748 $<0.001$**
Age -0.008 0.001 -13.845 $<0.001$**
Education -0.018 0.007 -2.542 0.011 *
Cond.(Forced):Veracity(True) -0.022 0.024 -0.928 0.354
@@
Cond.(HumanFC):Veracity(True) 0.121 0.034 3.554 $<0.001$**
Cond.(Forced):Congr.(Inc.) 0.015 0.016 0.988 0.323
Cond.(Optional):Congr.(Inc.) 0.051 0.013 3.982 $<0.001$**
Cond.(HumanFC):Congr.(Inc.) 0.056 0.028 1.970 0.049 *
@@
Cond.(Forced):Veracity(True):Congr.(Inc.) 0.031 0.022 1.368 0.171
Cond.(Optional):Veracity(True):Congr.(Inc.) -0.002 0.012 -0.186 0.852
Cond.(HumanFC):Veracity(True):Congr.(Inc.) -0.056 0.037 -1.504 0.133
@@
Figure S5 illustrates the relationship between belief in headlines and their congruency across all fact-checking scenarios and experimental conditions. The same relationship is presented with respect to sharing intent in Figure S6. We model this relationship using a three-way interaction between condition, fact-checking scenario, and headline congruence (Condition $\times$ FC Scenario $\times$ Congruence). Again, we focus on the forced and control conditions and exclude data for the optional participants when fitting each model. The results of fitting the belief and share group models are found in Tables S27 and S28. However, we again must utilize these models for post-hoc comparisons similar to those presented in the main text for each group. To do this, we compare headline congruence fitted slopes between the Control and Forced groups. These results are shown in Tables S29 and S30 for the belief and share group, respectively. We found no evidence of significant interactions within the belief group. However, in the sharing group, some significant interactions were observed for a specific fact-checking scenario. Participants who were forced to view unsure LLM fact checks about politically incongruent true headlines (True $\times$ unsure) were more likely to report a willingness to share these headlines compared to participants in the control group who viewed similar headlines. This was true despite the fact that, within each group, the tendency was to report a willingness to share incongruent headlines less than congruent headlines (Control: $b=-0.10$; LLM-forced: $b=-0.03$).
@@
Cond.(Forced):FC Scen.(True $\times$ false) -0.190 0.051 -3.683 $<0.001$**
@@
FC Scen.(False $\times$ unsure):Congr.(Inc.) -0.102 0.040 -2.536 0.011 *
@@
Cond.(Forced):FC Scen.(False $\times$ unsure):Congr.(Inc.) 0.008 0.045 0.173 0.863
@@
Cond.(Forced):FC Scen.(True $\times$ true):Congr.(Inc.) 0.069 0.035 1.962 0.050 *
@@
Cond.(Forced):FC Scen.(False $\times$ unsure):Congr.(Inc.) -0.088 0.056 -1.571 0.116
@@
Cond.(Forced):FC Scen.(True $\times$ true):Congr.(Inc.) -0.064 0.075 -0.855 0.393
@@
Headline Scenario Forced -- Control Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false 0.017 0.019 18738 0.873 1.000
False $\times$ unsure 0.025 0.058 18738 0.425 1.000
True $\times$ false 0.078 0.041 18738 1.887 0.296
True $\times$ true 0.086 0.048 18738 1.804 0.356
True $\times$ unsure 0.012 0.023 18738 0.511 1.000
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@
Headline Scenario Forced -- Control Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ false 0.025 0.020 19938 1.278 1.000
False $\times$ unsure -0.062 0.059 19938 -1.055 1.000
True $\times$ false 0.038 0.042 19938 0.919 1.000
True $\times$ true -0.039 0.048 19938 -0.815 1.000
True $\times$ unsure 0.071 0.023 19938 3.085 0.010 *
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
Next, we examine whether behaviors in the optional condition differ based on the congruence of headlines by introducing a three-way interaction term involving whether a participant chose to view LLM fact checks (opt in vs. opt out), fact-checking scenario, and headline congruence (Opt-Condition $\times$ FC Scenario $\times$ Congruence). Tables S31 and S32 show the results of fitting these models for the belief and intent to share groups, respectively. We perform a post-hoc comparison of the belief (Table S33) and sharing (Table S34) group slopes, fit to the opt-in and opt-out conditions across different levels of congruence. The results of these post-hoc comparisons are shown in Tables S35 and S36, respectively.
@@
Table S31: Opt In versus Opt Out Coefficients (Congruency interaction, Belief Group; $F=146.91$, $R^{2}=0.24$, $P<0.001$)
@@
(Intercept) 0.637 0.063 10.175 $<0.001$**
Option(opt out) -0.217 0.046 -4.765 $<0.001$**
FC Scen.(False $\times$ unsure) 0.072 0.083 0.864 0.388
FC Scen.(True $\times$ false) 0.034 0.018 1.856 0.063 $\cdot$
FC Scen.(True $\times$ true) 0.379 0.034 11.154 $<0.001$**
FC Scen.(True $\times$ unsure) 0.344 0.047 7.365 $<0.001$**
Congr.(Inc.) -0.003 0.023 -0.108 0.914
Age -0.004 0.001 -3.926 $<0.001$**
Education -0.009 0.009 -0.929 0.353
Option(opt out):FC Scen.(False $\times$ unsure) 0.068 0.096 0.710 0.478
Option(opt out):FC Scen.(True $\times$ false) 0.457 0.044 10.325 $<0.001$**
Option(opt out):FC Scen.(True $\times$ true) 0.193 0.049 3.907 $<0.001$**
Option(opt out):FC Scen.(True $\times$ unsure) 0.191 0.055 3.444 0.001 ***
Option(opt out):Congr.(Inc.) -0.054 0.043 -1.271 0.204
FC Scen.(False $\times$ unsure):Congr.(Inc.) 0.027 0.053 0.503 0.615
FC Scen.(True $\times$ false):Congr.(Inc.) 0.043 0.017 2.500 0.012 *
FC Scen.(True $\times$ true):Congr.(Inc.) 0.072 0.038 1.871 0.061 $\cdot$
FC Scen.(True $\times$ unsure):Congr.(Inc.) -0.083 0.039 -2.153 0.031 *
Option(opt out):FC Scen.(False $\times$ unsure):Congr.(Inc.) -0.165 0.073 -2.266 0.023 *
Option(opt out):FC Scen.(True $\times$ false):Congr.(Inc.) -0.032 0.052 -0.605 0.545
Option(opt out):FC Scen.(True $\times$ true):Congr.(Inc.) 0.024 0.064 0.382 0.703
Option(opt out):FC Scen.(True $\times$ unsure):Congr.(Inc.) 0.042 0.060 0.703 0.482
@@
Table S32: Opt In versus Opt Out Coefficients (Congruency interaction, Share Group; $F=113.92$, $R^{2}=0.19$, $P<0.001$)
@@
(Intercept) 0.798 0.068 11.778 $<0.001$**
Option(opt out) -0.304 0.039 -7.822 $<0.001$**
FC Scen.(False $\times$ unsure) 0.007 0.018 0.396 0.692
FC Scen.(True $\times$ false) -0.005 0.010 -0.462 0.644
FC Scen.(True $\times$ true) 0.084 0.039 2.179 0.029 *
FC Scen.(True $\times$ unsure) 0.098 0.023 4.253 $<0.001$**
Congr.(Inc.) 0.013 0.010 1.256 0.209
Age -0.006 0.001 -5.026 $<0.001$**
Education -0.001 0.014 -0.074 0.941
Option(opt out):FC Scen.(False $\times$ unsure) 0.043 0.029 1.472 0.141
Option(opt out):FC Scen.(True $\times$ false) 0.058 0.036 1.596 0.111
Option(opt out):FC Scen.(True $\times$ true) -0.008 0.068 -0.111 0.911
Option(opt out):FC Scen.(True $\times$ unsure) 0.030 0.034 0.897 0.370
Option(opt out):Congr.(Inc.) -0.022 0.014 -1.587 0.113
FC Scen.(False $\times$ unsure):Congr.(Inc.) 0.036 0.057 0.623 0.533
FC Scen.(True $\times$ false):Congr.(Inc.) -0.023 NaN
FC Scen.(True $\times$ true):Congr.(Inc.) 0.060 0.059 1.030 0.303
FC Scen.(True $\times$ unsure):Congr.(Inc.) -0.051 0.013 -3.937 $<0.001$**
Option(opt out):FC Scen.(False $\times$ unsure):Congr.(Inc.) -0.105 0.070 -1.494 0.135
Option(opt out):FC Scen.(True $\times$ false):Congr.(Inc.) 0.021 0.039 0.553 0.580
Option(opt out):FC Scen.(True $\times$ true):Congr.(Inc.) -0.069 0.097 -0.713 0.476
Option(opt out):FC Scen.(True $\times$ unsure):Congr.(Inc.) -0.028 0.030 -0.941 0.347
@@
Option Headline Scenario $b$ Std. Err. df $t$-ratio $P$ Sig.
Opt in False $\times$ false -0.003 0.019 9978 -0.136 0.891
Opt out False $\times$ false -0.057 0.018 9978 -3.077 0.002 **
Opt in False $\times$ unsure 0.024 0.052 9978 0.465 0.642
Opt out False $\times$ unsure -0.195 0.061 9978 -3.208 0.001 **
Opt in True $\times$ false 0.040 0.037 9978 1.081 0.280
Opt out True $\times$ false -0.046 0.042 9978 -1.098 0.272
Opt in True $\times$ true 0.069 0.043 9978 1.617 0.106
Opt out True $\times$ true 0.039 0.049 9978 0.813 0.416
Opt in True $\times$ unsure -0.086 0.020 9978 -4.257 $<$ .001 ***
Opt out True $\times$ unsure -0.097 0.024 9978 -4.085 $<$ .001 ***
@@
Option Headline Scenario $b$ Std. Err. df $t$-ratio $P$ Sig.
Opt in False $\times$ false 0.013 0.017 10138 0.729 0.466
Opt out False $\times$ false -0.010 0.020 10138 -0.480 0.632
Opt in False $\times$ unsure 0.048 0.052 10138 0.933 0.351
Opt out False $\times$ unsure -0.079 0.063 10138 -1.258 0.209
Opt in True $\times$ false -0.010 0.035 10138 -0.300 0.765
Opt out True $\times$ false -0.011 0.045 10138 -0.251 0.802
Opt in True $\times$ true 0.073 0.042 10138 1.758 0.079
Opt out True $\times$ true -0.018 0.052 10138 -0.352 0.725
Opt in True $\times$ unsure -0.038 0.019 10138 -1.961 0.050 *
Opt out True $\times$ unsure -0.088 0.025 10138 -3.501 $<$ 0.001 ***
@@
Headline scenario Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
True $\times$ False 0.086 0.056 9978 1.540 0.619
True $\times$ Unsure 0.012 0.031 9978 0.371 1.000
True $\times$ True 0.030 0.065 9978 0.459 1.000
False $\times$ False 0.054 0.026 9978 2.071 0.192
False $\times$ Unsure 0.219 0.080 9978 2.745 0.030 *
@@
$\dagger$ Bonferroni’s method comparing a family of 5 estimates
@@
Headline scenario Opt in -- Opt out Std. Error df $t$ ratio Adj. $P^{\dagger}$ Sig.
False $\times$ False 0.022 0.026 10138 0.838 1.000
False $\times$ Unsure 0.127 0.081 10138 1.564 0.589
True $\times$ False 0.001 0.057 10138 0.015 1.000
True $\times$ True 0.091 0.066 10138 1.375 0.845
True $\times$ Unsure 0.050 0.032 10138 1.577 0.574